OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. A(5, 5), m = 3

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

so u have x = 5 , y =5 and m is 3 , use the y = mx + c to find contant !!

OpenStudy (anonymous):

that's the standard line formula

OpenStudy (anonymous):

ok so how would it come out in the the formula like what would it look like

OpenStudy (anonymous):

come on , just try it !!

OpenStudy (anonymous):

do i plug it in

OpenStudy (anonymous):

tell me what u got and i'll check

OpenStudy (anonymous):

yes , plug it!!

OpenStudy (anonymous):

and then when i plug it in am i useing the like equation form to come up with the answer

OpenStudy (anonymous):

yes , exactly , except u only show the value of m & c ,

OpenStudy (anonymous):

so itll be like 5=m(5)+3

OpenStudy (anonymous):

no , m = 3 , use y = mx + c , to find c 5 = 3*5 + c , solve it

OpenStudy (anonymous):

then put the value back in the eqn. with leaving x & y as variables

OpenStudy (anonymous):

m=3 then you plug in the coordinates to x and y to find the b or y-intercept of the equation. Use the equation y=mx+b to find this. Then you have to turn that into standard form instead of point slope form.

OpenStudy (anonymous):

so plug in the what to the coordinates

OpenStudy (anonymous):

umm , u need the constant , first @ heat1017

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

use my eqn to get c then , i'll guide u

OpenStudy (anonymous):

5 = 5*3 + c 5 = 15 + c c = 5-15 = -10 then y = mx + c y = 3x -10

OpenStudy (anonymous):

What constant?

OpenStudy (anonymous):

But standard form of the answer is 3x+y=-10

OpenStudy (anonymous):

no , but then m = -3

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

then substitute -3 in the same eqn , u get ur answer

OpenStudy (anonymous):

Standard form Ax+By=b Point-Slope Form y=mx+b

OpenStudy (anonymous):

ok so what would be the answer then